OpenStudy (anonymous):

a softball players batting average is the ratio of hits the player makes to the number of times the player is at bat. if annas batting average is 0.320, how many hits would she expect to get in 25 times at bat?

OpenStudy (anonymous):

Let A = average, let b be times at bat, let h be hits A = h/b -> bA = h -> (25)(0.320) = h

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

sooo multplying?

OpenStudy (anonymous):

Yes, that's all you have to do. But it is good to see that by the formula and the re-write of the formula.

OpenStudy (anonymous):

okay thank you so much:D

OpenStudy (anonymous):

You could also try to do this in your head by the following, but it is a "backwards" approach: 0.320 = 320/1000 = 32/100 = 8/25 but that is a lot of mental gymnastics.

OpenStudy (anonymous):

Good luck in all of your studies, and thanks for the recognition! And you're quite welcome! @iloveGod